OpenStudy (anonymous):

A uniform linear charge of linear density y= 3.8nC/m, and it extends from x=0 to x=5, what is the total charge in nC?

OpenStudy (ash2326):

It extends from x=0 to x=5 so it extends for a length of 5 meters (5-0) Total Charge= \(y\times length\) \[=3.8 \ \frac{nC}{m}\times 5 m=?\] I think you can solve from here